OpenStudy (anonymous):

find sin(x-y) if sin x=3/5 and tan y = 7/24

OpenStudy (anonymous):

What's the formula for sin(x-y)?

OpenStudy (anonymous):

sin(x)cos(y)-cos(x)sin(y). I don't know how to put the tan in

OpenStudy (anonymous):

How is tan defined?

OpenStudy (anonymous):

sin/cos

OpenStudy (anonymous):

Can you relate sin to cos another way?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

assuming x and y are acute.

OpenStudy (anonymous):



OpenStudy (anonymous):

answer will be 44/125

OpenStudy (anonymous):

how come?

OpenStudy (anonymous):

sinx=3/5 tany=7/24 assuming x and y to b acute, u will get cos x=4/5, sin y=7/25, cos y= 24/25 put these values in the formula

OpenStudy (anonymous):

u will get the answer

OpenStudy (anonymous):

how did you get cos x=4/5?

OpenStudy (anonymous):

Did my diagram tell you anything about how sin and cos are related?

OpenStudy (anonymous):

hmm... \[\sin^2 +\cos^2=1 \] ?

OpenStudy (anonymous):

Yep. So what's cos x if sin x = 3/5?

OpenStudy (anonymous):

\[1-(3/5)^2\]

OpenStudy (anonymous):

The square root of that. Which is?

OpenStudy (anonymous):

AH! 4/5 thank you!

OpenStudy (anonymous):

similarly u can get sin y and cos y from the value of tan y

OpenStudy (anonymous):

Now, about the y. tan y = 7/24. So how would you get sin y and cos y from that?

OpenStudy (anonymous):

hmm...

OpenStudy (anonymous):

Remember that sines and cosines are legs of right triangles.

OpenStudy (anonymous):

oh! sin y is just 7 and cos is 24 ?

OpenStudy (anonymous):

opps oh wait

OpenStudy (anonymous):

 Yes, except cos and sin are always between -1 and 1. You have to scale them by the hypotenuse.